OpenStudy (anonymous):

While on a business trip to France, Bob felt ill. The hotel manager gave him a thermometer which measured his temperature in degrees Celsius. The thermometer read 38 degrees Celsius. The formula, F = 9/5C + 32, is used to convert temperatures in degrees (C) to degrees Fahrenheit (F). What is Bob's temperature in degrees Fahrenheit? Round the answer to the nearest whole number.

OpenStudy (anonymous):

Substitute C = 38 in the formula so F = (9/5)(38) + 32 F = 68.4 + 32 F = 100